# Energy Reflection and Transmission of Interfaces in $T\bar{T}$-deformed CFT

### v1

@@ 1 Introduction
The study of Interface Conformal Field Theories (ICFTs) has attracted considerable attention in recent years, with applications ranging from condensed matter physics to string theory. While for certain questions an ICFT can be mapped to a boundary conformal field theory (BCFT) by the folding procedure, there is a growing appreciation for the fact that ICFTs have distinctive features compared to a generic BCFT. These features can be seen, for instance, in the study of the entanglement between two CFTs joined by an interface (characterized by the so-called $g$-factor and the effective central charge) Karch:2023evr; Karch:2024udk; Afxonidis:2024gne, and in the transmission of energy between the two CFTs Quella:2006de; Meineri:2019ycm. The latter will be the topic of interest for this paper.
@@
A very remarkable fact about conformal interfaces in 2D CFTs is that the transmission of energy across the interface is universal Meineri:2019ycm, in the sense that it is independent of the details of the excitations that carry the energy (under some genericity assumptions that we will spell out later). To each interface one can assign a number, the energy transmission coefficient, or equivalently the reflection coefficient (as there is no absorption at a conformal interface in 2D), which determines the fraction of energy transmitted or reflected at the interface between the two bulk CFTs. [^1] We use the term bulk CFT for the CFT away from the interface, not to be confused with the bulk spacetime of the holographic duality. Hopefully the context makes it clear. In the absence of universality, one would describe the transmission by something like an S-matrix, that would depend on the incoming state, and would contain a priori infinitely many parameters. The universality reduces all these parameters to a single one. It also implies that, in the energy-momentum tensor sector, one can consider the conformal interface condition, i.e. the continuity of the energy flux across the interface, as a linear condition for the components of the energy-momentum tensor.
@@
The energy transmission coefficient has been studied for interfaces of free CFTs Bachas:2001vj, rational CFTs Furuta:2025ahl; Makabe:2017ygy, and holographic CFTs Bachas:2020yxv; Baig:2022cnb; Bachas:2022etu; Liu:2025khw. Apart from these few cases, not much is known in general; for instance, we do not know what is the behavior of the transmission coefficient under an RG flow, or under fusion of two interfaces. Part of the difficulty lies in the fact that both of these questions require breaking the conformal symmetry preserved by the interface. We are thus led to consider the following question: how is the energy transmitted across a non-conformal interface? In which way is the universality broken?
@@
The question is too broad to be answered in full generality. By breaking conformal invariance we lose all the tools that make the study of CFT tractable, and we have to rely on some form of perturbation theory. The only exception, to the best of our knowledge, is the $T\bar{T}$ deformation Zamolodchikov:2004ce; Dubovsky:2012wk; Caselle:2013dra; Cavaglia:2016oda; Smirnov:2016lqw, which allows one to move away from the conformal point and retain the solvability of the undeformed theory. Despite having the standard pathologies of an irrelevant deformation, namely non-locality, lack of UV-completeness, etc., it has gained a significant amount of interest owing to its integrability, factorization property, and holographic control. These remarkable properties have led to extensive study in the deformed theory in the context of deformed spectrum, S-matrix, entanglement properties, thermal transport properties, etc. (see e.g., the reviews Jiang:2019epa; He:2025ppz and references therein). However, the energy transport properties across an interface gluing a pair of deformed theories are yet to be investigated in detail, which is the goal of this paper. [^2] We should mention that there have been other studies of boundaries and interfaces in $T\bar{T}$-deformed theory, see Wang:2024jem.
@@
To study such energy transport properties, we will primarily make use of the fact that the $T\bar{T}$ deformation can be described as a dynamical (i.e. state-dependent) coordinate change Guica:2019nzm. The effect of the coordinate change can be encoded in a flow equation for the expectation value of the energy-momentum tensor, which is all that we need to extract the information of the energy transport. Notice that after we break the conformal symmetry in the bulk CFT, it is not obvious that this determines uniquely the breaking on the interface: one could consider additional deformations localized to the interface; without the constraint of conformal invariance, it is also not clear what conditions should be imposed to characterize the interface. One of the main points of this paper is that the flow equation gives a prescription to uniquely lift an interface defined in the undeformed CFT to the deformed theory. It is conceivable that the deformed theory might admit also other interfaces not determined by this procedure; we will not investigate this question.
@@
Since the flow equation is state-dependent, as mentioned above, in order to carry out the procedure explicitly we have to choose a particular state. We found that there are two cases that we can analyze explicitly. One is the NESS (non-equilibrium steady-state) created by taking two BCFTs at different temperatures, joining them at some initial time, and looking at the state created at late times (in the absence of the interface, this would be simply a boosted thermal state, and its $T\bar{T}$-deformation was studied in Medenjak:2020bpe; Medenjak:2020ppv). The second is a “scattering” state made of small plane-wave excitations around the vacuum. This case can be analyzed in a perturbative expansion in the amplitude of the fluctuations.
@@
After computing the energy transmission for these two states using the flow method, as a consistency check of our results, we shall reproduce them from the holographic computations in the bulk dual of these states. The $T\bar{T}$ deformation is known to admit a relatively simple holographic dual description, in terms of a finite cutoff McGough:2016lol or modified boundary conditions Guica:2019nzm. In the case of pure 3D gravity the two prescriptions coincide, and it can be shown that they directly yield the flow equations Guica:2019nzm. By analogy with the calculations of energy transmission in holographic CFT models, where the interface is described by a brane in the bulk AdS space Bachas:2020yxv; Bachas:2021tnp, we shall also consider gluing a pair of deformed geometries across a thin brane, within the modified boundary condition picture. We shall show that the holographic results match precisely with the results obtained with the flow method. Notice that, this is a non-trivial check of the prescription, since the presence of a brane introduces a source of matter in the bulk.
@@
The main result of this paper is the precise realization of the modified energy transport properties induced by the deformation. We have to distinguish the two cases. In the case of the NESS, we found that the linear matching conditions of the stress-tensor vevs Meineri:2019ycm across the conformal interface are generalized to non-linear conditions, wherein the single transmission coefficient becomes a transmission function, which is an antisymmetric function of the incoming energy fluxes on the interface. This function contains an infinite number of non-linear transmission coefficients, which can be considered universal in the same sense as the CFT coefficient: the amount of reflected and transmitted energy is only a function of the incoming fluxes and not of the way the energy is distributed in different modes of the theory. However, within our techniques, we cannot show that the universality holds in the sense of the independence on the operator used to create the excitation. We cannot find a closed form for the transmission function, but we can evaluate it to arbitrarily high order in the incoming fluxes. In the case of the scattering state, the time dependence introduces additional complications that make us restrict our analysis to linear order. For the scattering of single frequency states, we can define a frequency-dependent transmission coefficient, which however takes physical values only in the regime $\omega^{2}{\color[rgb]{0,0,0}\leq}\frac{6}{c\mid\mu\mid}$. For states that contain waves of different frequencies, the matching can be expressed as a function of the time derivatives of the incoming fluxes. Finally, within both the methods, we observe that the gluing of two deformed CFTs with central charges $c_{L,R}$ and deformation parameters $\mu_{L,R}$ must obey the condition $c_{L}\mu_{L}=c_{R}\mu_{R}$, which we conjecture to remain true in general. This condition has a very natural interpretation in the finite cutoff holographic prescription: it simply says that the radial cutoff must match on the two sides.
@@ 2 The universality of energy transmission in 1+1 dimensions
In this section, we shall briefly summarize the universality of the reflection and transmission coefficients of energy across a conformal interface gluing a pair of two-dimensional CFTs, following Meineri:2019ycm. Consider an interface gluing two 1+1 dimensional CFTs with central charges $c_{L}$ and $c_{R}$ respectively. Using translational invariance we shall fix the interface at $x=0$. The interface can be thought of as the worldline of the junction of two semi-infinite quantum wires glued at $x=0$. The residual symmetries are then the ones that leave the interface invariant, namely, translation and special conformal transformation along time, and dilatation. In this setup, we want to measure the coefficients of reflection and transmission of energy associated with the interface, which are given by
@@
Now the conservation law $\mathcal{T}_{L(R)}+\mathcal{R}_{L(R)}=1$ (seeMeineri:2019ycm for a detailed proof of this relation starting from the definitions (6)-(9)) leaves only two of the coefficients independent, which we choose to be $\mathcal{T}_{L,R}$. In what follows, we provide a derivation of the transmission coefficients under the simplifying assumption that the operator creating the excitation is holomorphic. For the proof with a generic operator, we refer to Meineri:2019ycm.
To evaluate $\mathcal{T}_{L}$, we need to compute the three-point function [^3] We switch to euclidean signature for the derivation. $\left\langle O_{L}^{1}(z_{1})T_{R}(z)O_{L}^{2}(z_{2})\right\rangle_{I}$. Here $O_{L}^{1},O_{L}^{2}$ are local operators in CFTL which we shall initially choose to be linear combinations of quasi-primaries. Note that for states created using holomorphic operators in CFTL, there is no dependence on the position of the wavepacket along $\bar{z}$ and the $D\rightarrow\infty$ limit can be dropped. Now, we shall first consider the OPE expansion of the two $O_{L}$s. Then the three-point function is essentially given by the sum of two-point functions of holomorphic operators of CFTL with $T_{R}$. Now the two-point function of a holomorphic operator with a generic operator in the presence of the interface is given by Meineri:2019ycm
@@
A similar computation for $\mathcal{T}_{R}$ would have led to $\mathcal{T}_{R}=c_{LR}/{c_{R}}$. Thus, the coefficients are completely fixed by the central charges of the two theories and by $c_{LR}$ Quella:2006de and are completely insensitive to the local operators used to create the states as well as to the choice of the wave packets. Note that, in addition to the conservation law, the transmission coefficients additionally obey $c_{L}\mathcal{T}_{L}=c_{R}\mathcal{T}_{R}$, which is the condition of detailed balance. This leaves only one out the four coefficients (6)-(9) independent. Finally note that, since the kinematic factors under the integral drop out between the numerator and the denominator, the above analysis will go through even if we take derivatives of the three-point function w.r.t $z_{1},z_{2}$ or multiply with $z,z^{2}$. This tells us that the above results are true for generic local operators in CFTL and also for the measurement of charges associated with dilatation or special conformal transformation.
@@ 3 $T\bar{T}$ deformation: flow equations and mixed boundary condition
Given the fact that $T\bar{T}$ is a double-trace deformation where the deforming operator also depends on the source, it is natural to expect the holographic dictionary for the deformed theory to be modified. In fact, such deformations are known to lead to mixed boundary conditions. In the context of $T\bar{T}$ deformation, the deformed dictionary can be most efficiently worked out using the variational principle Guica:2019nzm, which we briefly outline here. Taking a variation of the defining relation (13) with respect to the metric and after some simple algebra, one arrives at,
@@
where in arriving at (16) we have used the relations $\partial_{\mu}\sqrt{\gamma}=\sqrt{\gamma}\Theta$, $\Theta T_{\alpha\beta}-T_{\alpha\gamma}T_{\beta}^{\gamma}=-\frac{1}{2}\gamma_{\alpha\beta}\mathcal{O}_{T\bar{T}}$, and $\partial_{\mu}\Theta=-T^{\alpha\beta}T_{\alpha\beta}$ (see Appendix A of Guica:2019nzm; Banerjee:2024wtl). By differentiating (15) twice and using (16) we arrive at the final form of the flow equations
@@
Note that these solutions are non-perturbative in the deformation parameter. It is also useful to know that as a consequence of the flow equations, the following trace relation holds: [^4] In Guica:2022gts it is argued that this relation should be valid also at the quantum level up to total derivative terms, which were found to vanish at linear order in the perturbation.
@@ 4.1 Non-equilibrium Steady State (NESS)
This shows that the transmission coefficient can be read off from the Stefan-Boltzmann coefficient. The holographic version of this computation was done in Bachas:2021tnp.
@@ 4.2 Scattering state
where $i=L,R$. Note that, in this case, we are considering the unfolded setup unlike the folded setup of NESS state. Since the vevs are coordinate-dependent in this case, the pseudoconformal map (28) is difficult to invert. So we shall carry out the analysis perturbatively in the amplitude and the parameter $\epsilon$ is introduced to keep track of the order. To mimic the gravitational scattering setup of Bachas:2020yxv, we particularly choose ${\cal A}_{L}=1\,,{\cal A}_{R}={\cal T}_{L}^{(0)}{\cal A}_{L}\,,{\cal B}_{L}={\cal R}_{L}^{(0)}{\cal A}_{L}\,,{\cal B}_{R}=0$. This corresponds to a wave of amplitude unity incident on the interface from the left CFT, with the amplitude of the reflected(transmitted) wave being proportional to $\mathcal{R}^{(0)}(\mathcal{T}^{(0)})$. Similarly, for the conformal mappings, we choose the ansatz

### v2

@@ 1 Introduction
The study of Interface Conformal Field Theories (ICFTs) has attracted considerable attention in recent years, with applications ranging from condensed matter physics to string theory. While for certain questions an ICFT can be mapped to a boundary conformal field theory (BCFT) by the folding procedure, there is a growing appreciation for the fact that ICFTs have distinctive features compared to a generic BCFT. These features can be seen, for instance, in the study of the entanglement between two CFTs joined by an interface (characterized by the so-called $g$-factor and the effective central charge) Karch et al. (2023, 2024); Afxonidis et al. (2025), and in the transmission of energy between the two CFTs Quella et al. (2007); Meineri et al. (2020). The latter will be the topic of interest for this paper.
@@
A very remarkable fact about conformal interfaces in 2D CFTs is that the transmission of energy across the interface is universal Meineri et al. (2020), in the sense that it is independent of the details of the excitations that carry the energy (under some genericity assumptions that we will spell out later). To each interface one can assign a number, the energy transmission coefficient, or equivalently the reflection coefficient (as there is no absorption at a conformal interface in 2D), which determines the fraction of energy transmitted or reflected at the interface between the two bulk CFTs. In the absence of universality, one would describe the transmission by something like an S-matrix, that would depend on the incoming state, and would contain a priori infinitely many parameters. The universality reduces all these parameters to a single one. It also implies that, in the energy-momentum tensor sector, one can consider the conformal interface condition, i.e. the continuity of the energy flux across the interface, as a linear condition for the components of the energy-momentum tensor.
@@
The energy transmission coefficient has been studied for interfaces of free CFTs Bachas et al. (2002), rational CFTs Furuta et al. (2025); Makabe and Watts (2017), and holographic CFTs Bachas et al. (2020); Baig and Karch (2022); Bachas et al. (2023); Liu and Wang (2025). Apart from these few cases, not much is known in general; for instance, we do not know what is the behavior of the transmission coefficient under an RG flow, or under fusion of two interfaces. Part of the difficulty lies in the fact that both of these questions require breaking the conformal symmetry preserved by the interface. We are thus led to consider the following question: how is the energy transmitted across a non-conformal interface? In which way is the universality broken?
@@
The question is too broad to be answered in full generality. By breaking conformal invariance we lose all the tools that make the study of CFT tractable, and we have to rely on some form of perturbation theory. The only exception, to the best of our knowledge, is the $T\bar{T}$ deformation Zamolodchikov (2004); Dubovsky et al. (2012); Caselle et al. (2013); Cavaglià et al. (2016); Smirnov and Zamolodchikov (2017), which allows one to move away from the conformal point and retain the solvability of the undeformed theory. Despite having the standard pathologies of an irrelevant deformation, namely non-locality, lack of UV-completeness, etc., it has gained a significant amount of interest owing to its integrability, factorization property, and holographic control. These remarkable properties have led to extensive study in the deformed theory in the context of deformed spectrum, S-matrix, entanglement properties, thermal transport properties, etc. (see e.g., the reviews Jiang (2021); He et al. (2025) and references therein). However, the energy transport properties across an interface gluing a pair of deformed theories are yet to be investigated in detail, which is the goal of this paper. [^1] We should mention that there have been other studies of boundaries and interfaces in $T\bar{T}$-deformed theory, see Brizio et al. (2024); Wang and Deng (2025); Dey (2025).
@@
To study such energy transport properties, we will primarily make use of the fact that the $T\bar{T}$ deformation can be described as a dynamical (i.e. state-dependent) coordinate change Guica and Monten (2021). The effect of the coordinate change can be encoded in a flow equation for the expectation value of the energy-momentum tensor, which is all that we need to extract the information of the energy transport. Notice that after we break the conformal symmetry in the bulk CFT, it is not obvious that this determines uniquely the breaking on the interface: one could consider additional deformations localized to the interface; without the constraint of conformal invariance, it is also not clear what conditions should be imposed to characterize the interface. One of the main points of this paper is that the flow equation gives a prescription to uniquely lift an interface defined in the undeformed CFT to the deformed theory. It is conceivable that the deformed theory might admit also other interfaces not determined by this procedure; we will not investigate this question.
@@
Since the flow equation is state-dependent, as mentioned above, in order to carry out the procedure explicitly we have to choose a particular state. We found that there are two cases that we can analyze explicitly. One is the NESS (non-equilibrium steady-state) created by taking two BCFTs at different temperatures, joining them at some initial time, and looking at the state created at late times (in the absence of the interface, this would be simply a boosted thermal state, and its $T\bar{T}$-deformation was studied in Medenjak et al. (2021b, a)). The second is a “scattering” state made of small plane-wave excitations around the vacuum. This case can be analyzed in a perturbative expansion in the amplitude of the fluctuations.
@@
After computing the energy transmission for these two states using the flow method, as a consistency check of our results, we shall reproduce them from the holographic computations in the bulk dual of these states. The $T\bar{T}$ deformation is known to admit a relatively simple holographic dual description, in terms of a finite cutoff McGough et al. (2018) or modified boundary conditions Guica and Monten (2021). In the case of pure 3D gravity the two prescriptions coincide, and it can be shown that they directly yield the flow equations Guica and Monten (2021). By analogy with the calculations of energy transmission in holographic CFT models, where the interface is described by a brane in the bulk AdS space Bachas et al. (2020, 2021), we shall also consider gluing a pair of deformed geometries across a thin brane, within the modified boundary condition picture. We shall show that the holographic results match precisely with the results obtained with the flow method. Notice that, this is a non-trivial check of the prescription, since the presence of a brane introduces a source of matter in the bulk.
@@
The main result of this paper is the precise realization of the modified energy transport properties induced by the deformation. We have to distinguish the two cases. In the case of the NESS, we found that the linear matching conditions of the stress-tensor vevs Meineri et al. (2020) across the conformal interface are generalized to non-linear conditions, wherein the single transmission coefficient becomes a transmission function, which is an antisymmetric function of the incoming energy fluxes on the interface. This function contains an infinite number of non-linear transmission coefficients, which can be considered universal in the same sense as the CFT coefficient: the amount of reflected and transmitted energy is only a function of the incoming fluxes and not of the way the energy is distributed in different modes of the theory. However, within our techniques, we cannot show that the universality holds in the sense of the independence on the operator used to create the excitation. We cannot find a closed form for the transmission function, but we can evaluate it to arbitrarily high order in the incoming fluxes. In the case of the scattering state, the time dependence introduces additional complications that make us restrict our analysis to linear order. For the scattering of single frequency states, we can define a frequency-dependent transmission coefficient, which however takes physical values only in the regime $\omega^{2}{\color[rgb]{0,0,0}\leq}\frac{6}{c\mid\mu\mid}$. For states that contain waves of different frequencies, the matching can be expressed as a function of the time derivatives of the incoming fluxes. Finally, within both the methods, we observe that the gluing of two deformed CFTs with central charges $c_{L,R}$ and deformation parameters $\mu_{L,R}$ must obey the condition $c_{L}\mu_{L}=c_{R}\mu_{R}$, which we conjecture to remain true in general. This condition has a very natural interpretation in the finite cutoff holographic prescription: it simply says that the radial cutoff must match on the two sides.
@@ 2 The universality of energy transmission in 1+1 dimensions
In this section, we shall briefly summarize the universality of the reflection and transmission coefficients of energy across a conformal interface gluing a pair of two-dimensional CFTs, following Meineri et al. (2020). Consider an interface gluing two 1+1 dimensional CFTs with central charges $c_{L}$ and $c_{R}$ respectively. Using translational invariance we shall fix the interface at $x=0$. The interface can be thought of as the worldline of the junction of two semi-infinite quantum wires glued at $x=0$. The residual symmetries are then the ones that leave the interface invariant, namely, translation and special conformal transformation along time, and dilatation. In this setup, we want to measure the coefficients of reflection and transmission of energy associated with the interface, which are given by
@@
Now the conservation law $\mathcal{T}_{L(R)}+\mathcal{R}_{L(R)}=1$ (seeMeineri et al. (2020) for a detailed proof of this relation starting from the definitions (6)-(9)) leaves only two of the coefficients independent, which we choose to be $\mathcal{T}_{L,R}$. In what follows, we provide a derivation of the transmission coefficients under the simplifying assumption that the operator creating the excitation is holomorphic. For the proof with a generic operator, we refer to Meineri et al. (2020).
To evaluate $\mathcal{T}_{L}$, we need to compute the three-point function [^2] We switch to euclidean signature for the derivation. $\left\langle O_{L}^{1}(z_{1})T_{R}(z)O_{L}^{2}(z_{2})\right\rangle_{I}$. Here $O_{L}^{1},O_{L}^{2}$ are local operators in CFTL which we shall initially choose to be linear combinations of quasi-primaries. Note that for states created using holomorphic operators in CFTL, there is no dependence on the position of the wavepacket along $\bar{z}$ and the $D\rightarrow\infty$ limit can be dropped. Now, we shall first consider the OPE expansion of the two $O_{L}$s. Then the three-point function is essentially given by the sum of two-point functions of holomorphic operators of CFTL with $T_{R}$. Now the two-point function of a holomorphic operator with a generic operator in the presence of the interface is given by Meineri et al. (2020)
@@
A similar computation for $\mathcal{T}_{R}$ would have led to $\mathcal{T}_{R}=c_{LR}/{c_{R}}$. Thus, the coefficients are completely fixed by the central charges of the two theories and by $c_{LR}$ Quella et al. (2007) and are completely insensitive to the local operators used to create the states as well as to the choice of the wave packets. Note that, in addition to the conservation law, the transmission coefficients additionally obey $c_{L}\mathcal{T}_{L}=c_{R}\mathcal{T}_{R}$, which is the condition of detailed balance. This leaves only one out the four coefficients (6)-(9) independent. Finally note that, since the kinematic factors under the integral drop out between the numerator and the denominator, the above analysis will go through even if we take derivatives of the three-point function w.r.t $z_{1},z_{2}$ or multiply with $z,z^{2}$. This tells us that the above results are true for generic local operators in CFTL and also for the measurement of charges associated with dilatation or special conformal transformation.
@@ 3 $T\bar{T}$ deformation: flow equations and mixed boundary condition
Given the fact that $T\bar{T}$ is a double-trace deformation where the deforming operator also depends on the source, it is natural to expect the holographic dictionary for the deformed theory to be modified. In fact, such deformations are known to lead to mixed boundary conditions. In the context of $T\bar{T}$ deformation, the deformed dictionary can be most efficiently worked out using the variational principle Guica and Monten (2021), which we briefly outline here. Taking a variation of the defining relation (13) with respect to the metric and after some simple algebra, one arrives at,
@@
where in arriving at (16) we have used the relations $\partial_{\mu}\sqrt{\gamma}=\sqrt{\gamma}\Theta$, $\Theta T_{\alpha\beta}-T_{\alpha\gamma}T_{\beta}^{\gamma}=-\frac{1}{2}\gamma_{\alpha\beta}\mathcal{O}_{T\bar{T}}$, and $\partial_{\mu}\Theta=-T^{\alpha\beta}T_{\alpha\beta}$ (see Appendix A of Guica and Monten (2021); Banerjee and Roy (2024)). By differentiating (15) twice and using (16) we arrive at the final form of the flow equations
@@
Note that these solutions are non-perturbative in the deformation parameter. It is also useful to know that as a consequence of the flow equations, the following trace relation holds: [^3] In Guica et al. (2022) it is argued that this relation should be valid also at the quantum level up to total derivative terms, which were found to vanish at linear order in the perturbation.
@@ 4.1 Non-equilibrium Steady State (NESS)
This shows that the transmission coefficient can be read off from the Stefan-Boltzmann coefficient. The holographic version of this computation was done in Bachas et al. (2021).
@@ 4.2 Scattering state
where $i=L,R$. Note that, in this case, we are considering the unfolded setup unlike the folded setup of NESS state. Since the vevs are coordinate-dependent in this case, the pseudoconformal map (28) is difficult to invert. So we shall carry out the analysis perturbatively in the amplitude and the parameter $\epsilon$ is introduced to keep track of the order. To mimic the gravitational scattering setup of Bachas et al. (2020), we particularly choose ${\cal A}_{L}=1\,,{\cal A}_{R}={\cal T}_{L}^{(0)}{\cal A}_{L}\,,{\cal B}_{L}={\cal R}_{L}^{(0)}{\cal A}_{L}\,,{\cal B}_{R}=0$. This corresponds to a wave of amplitude unity incident on the interface from the left CFT, with the amplitude of the reflected(transmitted) wave being proportional to $\mathcal{R}^{(0)}(\mathcal{T}^{(0)})$. Similarly, for the conformal mappings, we choose the ansatz